At 8 a.m. an airplane left an airport and flew in a northerly direction at an average speed of 200 miles per hour. At 10 a.m. a jet plane which can average 600 miles per hour was sent out to overtake the first plane. At what time were both of the planes the same distance from the airport?
Answer by lwsshak3(11628)

At 8 a.m. an airplane left an airport and flew in a northerly direction at an average speed of 200 miles per hour. At 10 a.m. a jet plane which can average 600 miles per hour was sent out to overtake the first plane. At what time were both of the planes the same distance from the airport?
**
let x=travel time of 1st plane
x-2=travel time of jet plane
distance=travel time*speed (same for both planes)
..
200x=600(x-2)
200x=600x-1200
400x=1200
x=3
x-2=1
travel time of 1st plane=3 hrs
travel time of jet plane=1hr
Both planes were the same distance from the airport at 11 a.m.